When heated, KClO3
decomposes into KCl
and O2.


2KClO3⟶2KCl+3O2


If this reaction produced 77.8 g KCl,
how many grams of O2
were produced?
1
Expert's answer
2019-10-07T07:21:49-0400

Solution.

n(KCl) = 77.8/74.5 = 1.044 mole = n(KClO3)

1.044/2 = x/3

x = n(O2) = 1.566 moles

m(O2) = 1.566 * 32 = 50.112 g


Answer. 50.112 g